# From Specificity to Generality: Revisiting Generalizable Artifacts in Detecting Face Deepfakes

### v1

@@ 3.4 Loss Function
In addition to FIA-USA, which aims to generate fake training samples from real training samples, we also propose Automatic Forgery aware Feature Selection (AFFS) and Region aware Contrastive Regularization (RCR) based on FIA-USA.
@@ 3.4.1 Automatic Forgery-aware Feature Selection
Before the model training begins, AFFS calculates the similarity between the feature map differences of each dimension of the real and fake faces in the each layer of the neural network and the mask $M$, sorts from high to low and selects the top $m_{i}$ dimensions of features of the $i$-th layer to construct a Feature Pyramid Network (FPN) [49]. It is worth noting that AFFS is only performed before the model training starts, and once completed, which dimensions of each layer’s features will be used to construct the FPN is fixed. AFFS enables adaptive feature selection, reduces feature redundancy, enhances feature representativeness and discriminability, and ensures effective Region-aware Contrastive Regularization.
Specifically, given a neural network $\phi$, we define the output of $i$-th layer as $f_{i}=\phi_{i}(I)\in\mathbb{R}^{W_{i}*H_{i}*C_{i}}$ where $I\in\mathbb{R}^{H*W*3}$ represents the input image, $W_{i}$ and $H_{i}$ refer to the size of the feature map, and $C_{i}$ represents the feature dimension of $i$-th layer. Before training begins, we select the top $m_{i}$ channels for the $i$-th layer by minimizing the following loss function.
@@
$\mathcal{F}$ represents normalization and scaling operations, $f_{i}^{k}\in\mathbb{R}^{W_{i}*H_{i}},k\in[0,C_{i}]$, represents the feature map of the $k$-th dimension of the $f_{i}$ , $R_{n}$ and $F_{n}$ are positive and negative samples generated by FIA-USA, $M_{n}$ is the corresponding mask. Finally, we use the features $f_{i}^{\prime}$ selected by AFFS to construct FPN, where $f_{i}^{\prime}\in\mathbb{R}^{W_{i}*H_{i}*m_{i}}$.
@@ 3.4.2 Region-aware Contrastive Regularization
Given the network $\phi$ and the Feature Pyramid Network (FPN) $P$, we first define a set of triplets $F$, $R$, $M$, where $F$ and $R\in\mathbb{R}^{H*W*3}$ represents the pseudo-fake face and real face, and $M\in\mathbb{R}^{H*W}$ represents the corresponding mask in FIA-USA used to generate the fake face. We denote the feature maps of real and fake faces as $H_{R}=P(\phi(R))$, $H_{F}=P(\phi(F))$. For $H_{R}$, the areas inside and outside $M$ are real, while for $H_{F}$, the areas inside $M$ are manipulated and the areas outside $M$ are real. We define the region within $M$ as $H_{R}^{in}$, $H_{F}^{in}$ and the region outside $M$ as $H_{R}^{out}$, $H_{F}^{out}$. It is noteworthy that, aligning with LAA-Net [55], we acknowledge that pixels on the blend boundary are influenced by both false pixels within $M$ and real pixels outside $M$, hence we discard pixel features on the blend boundary.
@@
where $\delta(\cdot)$ is the normalized cosine similarity between two features, as:
@@
where $f^{i}$ represents a pixel feature. $\tau$ is a temperature parameter.
@@
where $\mathcal{L}_{BCE}$ denotes the cross-entropy classification loss. $\mathcal{L}_{BCE},$ $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ are weighted by the hyperparameters $\lambda_{1}$ ,$\lambda_{2}$ and $\lambda_{3}$, respectively.
@@ 4.1 Settings
FWA [45] Xception CVPRW’18 0.790 0.668 0.740 0.638 0.714
CapsuleNet [56] Capsule ICASSP’19 0.791 0.747 0.684 0.657 0.720
CNN-Aug [30] ResNet CVPR’20 0.742 0.703 0.646 0.617 0.677
Face X-ray [44] HRNet CVPR’20 0.709 0.679 0.766 0.694 0.712
FFD [11] Xception CVPR’20 0.784 0.744 0.802 0.743 0.768
F3Net [60] Xception ECCV’20 0.777 0.798 0.702 0.735 0.749
SPSL [51] Xception CVPR’20 0.815 0.765 0.812 0.741 0.783
SRM [53] Xception CVPR’21 0.793 0.755 0.812 0.741 0.775
CORE [57] Xception CVPRW’22 0.780 0.743 0.802 0.734 0.764
Recce [2] Designed CVPR’22 0.768 0.732 0.812 0.734 0.761
SBI* [66] EfficientNet-B4 CVPR’22 - 0.813 0.774 0.799 -
UCF [84] Xception ICCV’23 0.779 0.753 0.807 0.759 0.774
ED* [1] ResNet-34 AAAI’24 0.818 0.864 - 0.851 -
ProDet† [6] EfficientNet-B4 NeurIPS’24 0.909 0.842 0.848 0.774 0.843
LSDA* [86] EfficientNet-B4 CVPR’24 0.867 0.830 0.880 0.815 0.848
Forensic-Adapter* [10] CLIP (ViT-B/16) CVPR’25 - 0.837 - 0.799 -
@@
Table 2: Cross-dataset Evaluations: Cross-dataset evaluations were conducted using the Frame-level AUC. All experiments were trained on the c23 version of FF++ and tested on other datasets. ‘*’ indicates the results are cited from [10] and ‘†’ indicates our reproduction results using the checkpoints provided by the authors, otherwise, the results are from DeepfakeBench [85].
@@
Datasets.
To evaluate the effectiveness of our proposed method, we conducted extensive experiments on multiple datasets, including both traditional and generative deepfake datasets.
Traditional Deepfake Datasets. FaceForensics++ (FF++) [63], DeepfakeDetection (DFD) [12], Deepfake Detection Challenge (DFDC) [15], preview version of DFDC (DFDCP) [14], and CelebDF (CDF) [47]. FF++ comprises 1,000 original videos and 4,000 fake videos forged by four manipulation methods, namely, Deepfakes(DF) [13], Face2Face(F2F) [72], FaceSwap(FS) [18], and NeuralTextures(NT) [73]. Consistent with the SBI [67], we trained only with real faces from FF++. FF++ offers three levels of compression: raw, lightly compressed (c23), and heavily compressed (c40). Due to constraints regarding storage and download capabilities, we selected the c23 version of FF++, including DFD. Although many previous studies have utilized the same dataset for both training and testing, the preprocessing and experimental configurations can differ, making fair comparisons difficult. Therefore, in addition to testing on the raw data of the aforementioned datasets, we also performed generalization assessments on the unified new benchmark for traditional deepfakes, (DeepfakeBench) [85].
Method Venues uniface e4s facedancer fsgan inswap simswap Avg.
RECCE [2] CVPR 2022 84.2 65.2 78.3 88.4 79.5 73.0 78.1
SBI [67] CVPR 2022 64.4 69.0 44.7 87.9 63.3 56.8 64.4
CORE [57] CVPRW 2022 81.7 63.4 71.7 91.1 79.4 69.3 76.1
IID [34] CVPR 2023 79.5 71.0 79.0 86.4 74.4 64.0 75.7
UCF [84] ICCV 2023 78.7 69.2 80.0 88.1 76.8 64.9 77.5
LSDA [86] CVPR 2024 85.4 68.4 75.9 83.2 81.0 72.7 77.8
CDFA [50] ECCV 2024 76.5 67.4 75.4 84.8 72.0 76.1 75.9
ProgressiveDet [6] NeurIPS 2024 84.5 71.0 73.6 86.5 78.8 77.8 78.7
Ours - 91.8 87.5 83.0 86.3 87.4 91.0 87.8
Table 3: We compared our method with state-of-the-art detection methods on six representative face swapping forgeries in DF40 [87].
Generative Deepfake Datasets.
Diffusion Facial Forgery (DiFF) [5] contains over 500,000 images synthesized using 13 different generation methods under four conditions: Text to Image (T2I), Image to Image (I2I), Face Swapping (FS) and Face Editing (FE). Text to Image encompasses four generation methods: Midjourney [54], Stable Diffusion XL (SDXL) [59], Free-DoM_T [88], and HPS [82]. Image to Image comprises Low Rank Adaptation(LoRA) [32], DreamBooth [64], SDXL Refiner [59], and Free-DoM_I. Face Swapping includes DiffFace [40] and DCFace [40]. Face Editing comprises Imagic [39], Cycle Diffusion(CycleDiff) [80], and Collaborative Diffusion(CoDiff) [35]. Besides, we also conducted experiments on the latest deepfake dataset DF40 [87], which contains the forgery data generated within the FF++ domain, ensuring the evaluation fake methods different but the data domain unchanged.
Evaluation Metrics.
We report the Frame-Level Area Under Curve (AUC) metric on the DeepfakeBench and DIFF dataset, report the Video-Level Area Under Curve (AUC) metric on the traditional deepfake datasets, to compare our proposed method with prior works.
Implementation Details.
We utilize EfficientNetB4 [71] as the default neural network, and employ the SAM [21] optimizer to train it for 50 epochs. The batch size is set to 12. We uniformly sample 32 frames from each video during training. The initial learning rate is 0.001. For data augmentation, in addition to the proposed data augmentation (FIA-USA), RandomHorizontalFlip, RandomCutOut, and AddGaussian Noise were also applied. All experiments were conducted on a single NVIDIA 3090. We also investigate alternative network architectures and their respective outcomes. Empirically, $\lambda_{1},\lambda_{2},\lambda_{3}$ are assigned values of 1, 2.5, 0.25. The temperature parameter $\tau$ is 0.7. We also explored the impact of other variants on the detection results.
4.2 Generalization Evaluation
Our model’s generalization was evaluated on traditional and generation deepfake datasets at both the frame and video levels.
4.2.1 Traditional Deepfake Datasets
Model Venues CDF-v2 DFDC DFDCP
Face X-Ray [44] CVPR’20 - - 0.711
LipForensics [28] CVPR’21 0.824 0.735 -
FTCN [94] ICCV’21 0.869 - 0.740
PCL + I2G [93] ICCV’21 0.900 0.675 0.743
HCIL [25] ECCV’22 0.790 0.692 -
ICT  [17] CVPR’22 0.857 - -
SBI [67] CVPR’22 0.928 0.719 0.855
AltFreezing [79] CVPR’23 0.895 - -
SAM [7] CVPR’24 0.890 - -
LSDA [86] CVPR’24 0.911 0.770 -
CFM [52] TIFS’24 0.897 - 0.802
Ours - 0.941 0.732 0.866
Table 4: Comparison with SOTA methods using the Video-Level AUC. The results of other methods are cited from their original papers.
Frame-level.
We first compare our method with previous work at the frame level. To ensure the fairness of the experiment, we conducted the experiment on the DeepfakeBench, and adhered to the data preprocessing and experimental settings provided by them. For previous work, we utilized the experiment data provided by DeepfakeBench. As shown in Table 2, our method outperforms other methods on the majority of deepfake datasets and competes with state-of-the-art methods on the CDF-v1 dataset.
Video-Level.
In addition to comparing at the Frame-Level, we also compared our method at the Video-Level with state-of-the-art detection algorithms, including various data augmentation methods. Typically, the experimental results are directly taken from the original paper and trained on the c23 version of FF++. FreqBlender [43] did not provide experimental results trained on the FF++ c23 version in their paper, therefore we directly cited experiment results on the FF++ raw version. The comparison results, as shown in Table 4, strongly demonstrate the effective generalization of our method, and our method outperforms other methods, even being comparable to FreqBlender trained on the raw version of FF++.
Method Train Set Test Subset
Real Fake T2I I2I FS FE
Xception [63] 62.43 56.83 85.97 58.64
F3-Net [60] 66.87 67.64 81.01 60.60
EfficientNet [71] 74.12 57.27 82.11 57.20
DIRE [78] 44.22 64.64 84.98 57.72
SBI [67] 80.20 80.40 85.08 68.79
Ours 86.05 84.95 89.42 72.73
Table 5: Comparison with universal deepfake detection methods using the Frame-Level AUC on the DiFF dataset. ‘†’ indicates models that are designed for deepfake detection, while ‘‡’ signifies models intended for general generated image detection.
@@ 4.2.1 Traditional Deepfake Datasets
(a) Ablation Study of Method Components. We report the Video-Level AUC for traditional deepfake datasets and the Frame-Level AUC for generative deepfake datasets. Mmr denotes multiple types of masks, SR signifies the process of super-resolution, and AE refers to the reconstruction achieved through autoencoders.
@@
CDF-v2 DFDCP T2I I2I FS FE
Res50 82.45 79.14 74.73 76.94 68.82 58.59 73.44
Res101 85.51 82.35 73.21 74.95 71.26 60.44 74.62
Effb1 91.88 82.30 83.86 85.24 83.73 70.69 82.86
Effb4 94.10 86.66 86.05 84.95 89.42 72.73 85.65
(b) Ablation Study on Model Architectures.
@@
4.2.2 Generation Deepfake Datasets
As mentioned in Section 2, researchers have found that RGB enhancement methods exhibit limited efficacy when facing generative deepfakes. Here, we compared multiple universal deepfake detection methods on the generative deepfake dataset DIFF, as shown in Table 3 and Table 5. Our method not only exhibits robust generalization capabilities on traditional deepfakes, but also showcases superior performance in generative deepfake detection. Compared to the baseline method, our approach achieved a significant improvement of nearly 5%.
In section G of the appendix, we conducted experiments on more forgery techniques provided by DF40 and DIFF.
4.2.3 Robustness
In Figure 3, we assessed the influence of different levels of random perturbations on the detection performance, the experimental setup adhering to the experimental protocols established in prior studies. We quantified the effects of different levels of Gaussian Blur, Block Wise, Change Contrast, Change Saturation, and JPEG Compression on the detectors. Notably, to gauge accuracy, we employed the reduction rate of Video-Level AUC to assess the robustness of the detector to different degrees of disturbance, which could be expressed as $R_{AUC}=(AUC-AUC_{raw})/AUC_{raw}$, where $AUC_{raw}$ refers to no perturbations. Figure 3 shows that our method is more robust relative to other RGB-based methods. Change Contrast, Change Saturation, and JPEG Compression on the detectors. Notably, to gauge accuracy, we employed the reduction rate of Video-Level AUC to assess the robustness of the detector to different degrees of disturbance, which could be expressed as $R_{AUC}=(AUC-AUC_{raw})/AUC_{raw}$, where $AUC_{raw}$ refers to no perturbations. Figure 3 shows that our method is more robust relative to other RGB-based methods.

### v2

@@ 3.2.1 Automatic Forgery-aware Feature Selection
While FIA-USA generates discriminative training samples through dual-artifact augmentation, the inherent feature redundancy in Feature Pyramid Networks (FPN) lin2017feature may drive models to focus on non-critical forgery patterns, thereby compromising detection generalization. This limitation stems from FPN’s naive aggregation of all channel-wise features without adaptively emphasizing FIA/USA-correlated representations. To address this limitation, we propose Automatic Forgery-aware Feature Selection (AFFS), a statistically-driven feature compression paradigm that quantifies channel-wise sensitivity to forgery regions, thereby dynamically constructing lightweight yet discriminative feature pyramids.
Given a pretrained backbone network $\phi$, let $f_{i}=\phi_{i}(I)\in\mathbb{R}^{W_{i}*H_{i}*C_{i}}$ denote the feature map from the $i$-th layer, where $I\in\mathbb{R}^{H*W*3}$ represents the input image. To identify artifact-sensitive feature dimensions, we leverage pseudo-fake pairs $\{R_{n},F_{n},M_{n}\}_{n=1}^{N}$ generated by FIA-USA and compute the normalized response discrepancy between real and forged regions. For the $k$-th channel $f_{i}^{k}\in\mathbb{R}^{W_{i}*H_{i}}$ in layer $i$, the selection criterion is formulated as:
@@
where $\mathcal{F}$ denotes feature normalization and spatial interpolation to align $f_{i}^{k}$ with mask $M_{n}$. Channels are ranked by ascending $\mathcal{L}_{AFFS}^{i,k}$ and the top-$m_{i}$ channels with minimal errors are retained to form a compressed feature map $f_{i}^{\prime}\in\mathbb{R}^{W_{i}*H_{i}*m_{i}}$ for FPN construction. AFFS leverages a pre-trained model as initialization, operating as a supervised feature ranking mechanism that reduces feature redundancy.
@@ 3.2.2 Region-aware Contrastive Regularization
Building upon the artifact-sensitive features selected by AFFS, we design Region-aware Contrastive Regularization (RCR) to explicitly model the divergence between forged and authentic regions through multi-granularity contrastive learning. Given the Feature Pyramid Network (FPN) $P$, we denote the feature maps of real and fake faces as $H_{R}=P(\phi(R))$, $H_{F}=P(\phi(F))$. For $H_{R}$, the areas inside and outside $M$ are real, while for $H_{F}$, the areas inside $M$ are manipulated and the areas outside $M$ are real. We define the region within $M$ as $H_{R}^{in}$, $H_{F}^{in}$ and the region outside $M$ as $H_{R}^{out}$, $H_{F}^{out}$. RCR operates on the refined feature pyramid from AFFS, implementing multi-granularity contrast through:
@@
where $f^{i}$ represents a pixel feature, $\tau$ is a temperature parameter and $\delta(\cdot)$ is the normalized cosine similarity between two features, as:
@@
where $\mathcal{L}_{BCE}$ denotes the cross-entropy classification loss. $\mathcal{L}_{BCE},$ $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ are weighted by the hyperparameters $\lambda_{1}$ $\lambda_{2}$ and $\lambda_{3}$, respectively.
@@ 4.1 Settings
Datasets. To evaluate the effectiveness of our proposed method, we conducted extensive experiments on seven widely-adopted benchmark datasets spanning both classical facial manipulation paradigms and emerging generative deepfake architectures.a. Traditional Deepfake Datasets: 1. FaceForensics++ (FF++) rossler2019faceforensics++ , 2. Deepfake Detection (DFD) dfd , 3. Deepfake Detection Challenge (DFDC) dolhansky2020deepfake , 4. preview version of DFDC (DFDCP) dolhansky2019dee , and 5. CelebDF (CDF) li2020celeb . FF++ comprises 1,000 original videos and 4,000 fake videos forged by four manipulation methods, namely, Deepfakes(DF) df , Face2Face(F2F) thies2016face2face , FaceSwap(FS) fs , and NeuralTextures(NT) thies2019deferred . FF++ offers three levels of compression: raw, lightly compressed (c23), and heavily compressed (c40), under storage constraints, our implementation adopts the FF++_c23 (including DFD) with training conducted following the SBI protocol shiohara2022detecting , employing exclusively real facial samples from FF++_c23 subset. Although many previous studies have utilized the same dataset for both training and testing, the preprocessing and experimental configurations can differ, making fair comparisons difficult. Therefore, in addition to testing on the raw data of the aforementioned datasets, we also performed generalization assessments on the unified new benchmark for traditional deepfakes, (DeepfakeBench) yan2023deepfakebench . b. Generative Deepfake Datasets. 6. Diffusion Facial Forgery (DiFF) cheng2024diffusion contains over 500,000 images synthesized using 13 different generation methods under four conditions: Text to Image (T2I), Image to Image (I2I), Face Swapping (FS) and Face Editing (FE). Text to Image encompasses four generation methods: Midjourney Midjourney , Stable Diffusion XL (SDXL) podell2023sdxl , Free-DoM_T yu2023freedom , and HPS wu2023better . Image to Image comprises Low Rank Adaptation(LoRA) hu2021lora , DreamBooth DreamBooth , SDXL Refiner podell2023sdxl , and Free-DoM_I. Face Swapping includes DiffFace kim2022diffface and DCFace kim2022diffface . Face Editing comprises Imagic kawar2023imagic , Cycle Diffusion(CycleDiff) wu2023latent , and Collaborative Diffusion(CoDiff) huang2023collaborative . 7. DF40 yan2024df40 encompasses 40 state-of-the-art deepfake generation methodologies spanning four core categories: face swapping, facial reenactment, full-face synthesis, and advanced facial manipulation. Significantly surpassing existing benchmarks in both scope and volume, DF40 integrates cutting-edge generative architectures from 2024 alongside widely adopted commercial software solutions.
Evaluation Metrics. We report the Frame-Level Area Under Curve (AUC) metric on the DeepfakeBench and DIFF dataset, report the Video-Level Area Under Curve (AUC) metric on the traditional deepfake datasets, to compare our proposed method with prior works.
Implementation Details. We adopt EfficientNetB4 tan2019efficientnet as the backbone network architecture (We also investigate alternative network architectures and their respective outcomes), trained for 50 epochs using the SAM optimizer foret2020sharpness with a batch size of 12 and initial learning rate of 0.001. For video processing, each input is uniformly sampled into 32 frames during both training and inference phases. Our data augmentation pipeline combines the proposed FIA-USA strategy with conventional techniques including RandomHorizontalFlip, RandomCutOut, and AddGaussianNoise. The loss coefficients $\lambda_{1},\lambda_{2},\lambda_{3}$ are empirically set to 1, 2.5, and 0.25 respectively (We also explored the impact of other variants on the detection results), with the temperature parameter $\tau$ fixed at 0.7. All experiments were conducted on a single NVIDIA 3090.
Table 2: Cross-dataset Evaluations: Cross-dataset evaluations were conducted using the Frame-level AUC. All experiments were trained on the c23 version of FF++ and tested on other datasets. * indicates the results are cited from cui2024forensics and † indicates our reproduction results using the checkpoints provided by the authors, otherwise, the results are from DeepfakeBench yan2023deepfakebench .
@@
FWA li2018exposing Xception CVPRW’18 0.790 0.668 0.740 0.638 0.714
CapsuleNet nguyen2019capsule Capsule ICASSP’19 0.791 0.747 0.684 0.657 0.720
CNN-Aug he2016deep ResNet CVPR’20 0.742 0.703 0.646 0.617 0.677
Face X-ray li2020face HRNet CVPR’20 0.709 0.679 0.766 0.694 0.712
FFD dang2020detection Xception CVPR’20 0.784 0.744 0.802 0.743 0.768
F3Net qian2020thinking Xception ECCV’20 0.777 0.798 0.702 0.735 0.749
SPSL liu2021spatial Xception CVPR’20 0.815 0.765 0.812 0.741 0.783
SRM luo2021generalizing Xception CVPR’21 0.793 0.755 0.812 0.741 0.775
CORE ni2022core Xception CVPRW’22 0.780 0.743 0.802 0.734 0.764
Recce cao2022end Designed CVPR’22 0.768 0.732 0.812 0.734 0.761
SBI* shao2022detecting EfficientNet-B4 CVPR’22 - 0.813 0.774 0.799 -
UCF yan2023ucf Xception ICCV’23 0.779 0.753 0.807 0.759 0.774
ED* ba2024exposing ResNet-34 AAAI’24 0.818 0.864 - 0.851 -
ProDet† cheng2024can EfficientNet-B4 NeurIPS’24 0.909 0.842 0.848 0.774 0.843
LSDA* yan2024transcending EfficientNet-B4 CVPR’24 0.867 0.830 0.880 0.815 0.848
Forensic-Adapter* cui2024forensics CLIP (ViT-B/16) CVPR’25 - 0.837 - 0.799 -
@@
4.2 Generalization Evaluation
4.2.1 Traditional Deepfake Datasets.
We first compare our method with previous work at the frame level. To ensure the fairness of the experiment, we conducted the experiment on the DeepfakeBench, and adhered to the data preprocessing and experimental settings provided by them. For previous work, we utilized the experiment data provided by DeepfakeBench. As shown in Table 2, our method outperforms other methods on the majority of deepfake datasets and competes with state-of-the-art methods on the CDF-v1 dataset. In addition to comparing at the frame-level, we also compared our method at the video-level with state-of-the-art detection algorithms, including various data augmentation methods. The comparison results, as shown in Table 3, strongly demonstrate the effective generalization of our method. Table 3: Comparison with SOTA methods using the Video-Level AUC. Model Venues CDF-v2 DFDC DFDCP Face X-Ray li2020face CVPR’20 - - 0.711 LipForensics haliassos2021lips CVPR’21 0.824 0.735 - FTCN zheng2021exploring ICCV’21 0.869 - 0.740 PCL + I2G zhao2021learning ICCV’21 0.900 0.675 0.743 HCIL gu2022hierarchical ECCV’22 0.790 0.692 - ICT  dong2022protecting CVPR’22 0.857 - - SBI shiohara2022detecting CVPR’22 0.928 0.719 0.855 AltFreezing wang2023altfreezing CVPR’23 0.895 - - SAM Choi_2024_CVPR CVPR’24 0.890 - - LSDA yan2024transcending CVPR’24 0.911 0.770 - CFM luocfm TIFS’24 0.897 - 0.802 LAA-Net† nguyen2024laa CVPR’24 0.840 - 0.741 Ours - 0.941 0.732 0.866 Table 4: Comparison with universal deepfake detection methods using the Frame-Level AUC on the DiFF dataset. $\dagger$} indicates models that are designed for deepfake detection, while $\ddagger$} signifies models intended for generated image detection. Method Test Subset T2I I2I FS FE Xception rossler2019faceforensics++ 62.43 56.83 85.97 58.64 F3-Net qian2020thinking 66.87 67.64 81.01 60.60 EfficientNet tan2019efficientnet 74.12 57.27 82.11 57.20 DIRE wang2023dire 44.22 64.64 84.98 57.72 SBI shiohara2022detecting 80.20 80.40 85.08 68.79 Ours 86.05 84.95 89.42 72.73
Table 5: Comparative Analysis of Detection Performance Between the Proposed Method and State-of-the-Art Approaches on Six Representative Face Swapping Forgery Types in DF40 yan2024df40 .
Method Venues uniface e4s facedancer fsgan inswap simswap Avg.
RECCE cao2022end CVPR 2022 84.2 65.2 78.3 88.4 79.5 73.0 78.1
SBI shiohara2022detecting CVPR 2022 64.4 69.0 44.7 87.9 63.3 56.8 64.4
CORE ni2022core CVPRW 2022 81.7 63.4 71.7 91.1 79.4 69.3 76.1
IID huang2023implicit CVPR 2023 79.5 71.0 79.0 86.4 74.4 64.0 75.7
UCF yan2023ucf ICCV 2023 78.7 69.2 80.0 88.1 76.8 64.9 77.5
LSDA yan2024transcending CVPR 2024 85.4 68.4 75.9 83.2 81.0 72.7 77.8
CDFA lin2024fake ECCV 2024 76.5 67.4 75.4 84.8 72.0 76.1 75.9
ProgressiveDet cheng2024can NeurIPS 2024 84.5 71.0 73.6 86.5 78.8 77.8 78.7
Ours - 91.8 87.5 83.0 86.3 87.4 91.0 87.8
4.2.2 Generative Deepfake Datasets
Our evaluation on DIFF and DF40 datasets demonstrates the superior generalization capabilities of our architecture compared to state-of-the-art detection methods. Comprehensive metrics are detailed in Table 4 and Table 5. As shown in Table 5, our method achieves an average AUC of 87.8%, surpassing RECCE by 9.7% and outperforming SBI by 23.4%, and enables 10% higher AUC than LSDA’s latent space augmentation on average. The results notably contradict recent findings yan2024transcending about RGB-based methods’ limitations against generative deepfakes, proving that properly designed RGB-level artifacts retain critical discriminative signals. For comprehensive experimental results on forgery techniques supported by the DF40 and DIFF, please refer to the extended analysis in Appendix Section G.
4.2.3 Robustness
In Figure 3, we assessed the influence of different levels of random perturbations on the detection performance. We quantified the effects of different levels of Gaussian Blur, Block Wise, Change Contrast, and JPEG Compression on the detectors. Notably, to gauge accuracy, we employed the reduction rate of Video-Level AUC to assess the robustness of the detector to different degrees of disturbance, which could be expressed as $R_{AUC}=(AUC-AUC_{raw})/AUC_{raw}$, where $AUC_{raw}$ refers to no perturbations. As shown in the Figure 3, our method exhibits superior robustness compared to other methods.
@@ 4.2.3 Robustness
(a) Ablation Study of Method Components. We report the Video-Level AUC for traditional deepfake datasets and the Frame-Level AUC for generative deepfake datasets. SR signifies the process of super-resolution, and AE refers to the reconstruction achieved through autoencoders.
@@
(b) Ablation Study on Model Architectures.
@@
CDF-v2 DFDCP FS FE
Res50 82.45 79.14 68.82 58.59 72.25
Res101 85.51 82.35 71.26 60.44 74.89
Effb1 91.88 82.30 83.73 70.69 82.15
Effb4 94.10 86.66 89.42 72.73 85.72
(c) Experiment on Hyperparameter Configuration.